OpenStudy (anonymous):

Solve the logarithmic equation. Be sure to reject any value that is not in the domain of the original logarithmic expressions. Give the exact answer. log6x2 = log6(5x + 36)

OpenStudy (anonymous):

choices A. (3/2) B. {9} C. Ø D. {9, -4}

OpenStudy (anonymous):

@Vaidehi09

OpenStudy (anonymous):

simplify the eq: log 6 + logx^2 = log6 + log(5x+36) logx^2 = log(5x+36) taking antilogs, x^2 = 5x + 36 now, this is a quadratic eq. u should know ur way now.

OpenStudy (anonymous):

thnxx!